OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 12(1.03)n Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent? (4 points)

OpenStudy (mathstudent55):

n is an exponent, right?

OpenStudy (anonymous):

yah

OpenStudy (mathstudent55):

You mean this function: \(\large f(n) = 12(1.03)^n\)

OpenStudy (mathstudent55):

Set the function equal to 16.3 and solve for n. That will give you an idea of the domain you need for n.

OpenStudy (anonymous):

how do I do that?

OpenStudy (mathstudent55):

You need to use logs to solve for an exponent.

OpenStudy (anonymous):

idkhow

OpenStudy (mathstudent55):

Here is an example: \(24 = 4(5.5)^x\) First, switch sides to have the variable on the left side. \(4(5.5)^x = 24\) Divide both sides by 4 \(\dfrac{4(5.5)^x}{4} =\dfrac{24}{4}\) \((5.5)^x=6\) Take log of both sides \(\log (5.5)^x = \log 6\)] The log of an exponent is a multiplier \(x \log 5.5 = \log 6\) Divide both sides by \(\log 5.5\) \(x = \dfrac{\log 6}{\log 5.5} \) \(x \approx \dfrac{0.77815}{0.74036}\) \(x \approx 1.0510\)